Hi,
I am trying to help my son with this applied math 10 problem and I need some help.
A car is travelling at 60 km/h West away from Vancouver. It is 200 km away from Vancouver after 3 hours. Write an equation which models this situation. Use the equation to determine the distance away from Vancouver after 8 hrs.
m = _______________ (x,y) = ( )
Could you help me please?
Thank you and Happy New Year,
Phill


Answer by stanbon(75887)

I am trying to help my son with this applied math 10 problem and I need some help.
A car is travelling at 60 km/h West away from Vancouver. It is 200 km away from Vancouver after 3 hours. Write an equation which models this situation. Use the equation to determine the distance away from Vancouver after 8 hrs.
m = _______________ (x,y) = ( )
=====================================
1st: Draw the picture
2nd: Equation:
distance = rate * time
d(t) = r*t
After t hours the car will have gone 60*t km
-------
d(3) = 60*3 = 180 km
---
d(8) = 60*8 = 480 km
============================
Cheers,
Stan H.